OpenStudy (anonymous):

find angle A if B=95, b=15.4 and a=8.0

OpenStudy (anonymous):

Use the Law of Sines. \[\frac{\sin(A)}{a}=\frac{\sin(B)}{b}\]

OpenStudy (anonymous):

ok i got sin(A)/8.0= sin(95)/15.4 what do i do now?

OpenStudy (anonymous):

??